The area of a rectangular field is
3864

yd2
. If the length of the field is
69
yards, what is its width?

Answer by checkley77(12844)

A=LW
3864=69*W
W=3864/69
W=56
ANS. FOR THE WIDTH.
PROOF:
3864=69*56
3864=3864